OpenStudy (anonymous):

Given a positive integer n, suppose S is a subset of {1, 2,..., 2n} with |S| = n + 1. Prove that there are distinct a,b in S such that a divides b

myininaya (myininaya):

What have you tried? Or what has been your thinking so far. Do you understand what it is asking or saying? Like so we have the universe={1,2,3,...,n,n+1,n+2,n+3,...n+n} (n+n=2n I just wrote it like the above to see the set better for myself) Now S is a subset of that universe and is said to have (n+1) elements of the universe. There are suppose to be numbers a and b in S such that a|b or a*(some integer)=b Now if n=1 we have the universe={1,2(1)}={1,2} There is exactly 2 elements in the universe if n=1 Since n=1, then |S|=1+1=2 Which means that S has to be the universe and 1 does divide 2 since 1(2)=2. If n=5 we have the universe={1,2,3,4,5,5+1,5+2,5+3,5+4,5+5} ={1,2,3,4,5,6,7,8,9,10} Remember |S|=1+5=6 So we are suppose to have 6 elements in the universe. There are many different sets that can be obtained for S. Here are some possibilities for S: Possibility 1: {1,2,3,4,5,6} 1|any number 2|4 2|6 3|6 Possibility 2: {1,2,3,4,5,7} 1| any number 2|4 Possibility 3: {1,2,3,4,5,10} 1|any number 2|4 2|10 Possibility 4: {1,2,3,7,9,10} 1|any number 2|10 3|9 Possibility 5: {2,3,5,7,9,10} 2|10 3|9 I will have to be back but I will do some more thinking. Be back in maybe like 3 hours. Sorry. Class time.